# Rethinking Feature Backbone Fine-tuning for Remote Sensing Object Detection

### v1

@@ IV Conclusions and Future Work
In this letter, we have proposed a simple but efficient feature backbone fine-tuning strategy named DBF for remote sensing object detection. This work showed that alternating between freezing and unfreezing the backbone aids object detection models in learning both general and domain-specific knowledge. Especially with the proposed method, a significant decrease in memory usage and training time can be achieved. In this work, our approach has shown powerful results on DOTA and DIOR-R with various detection architectures.
@@
In the future, more sophisticated studies will be conducted as follows: (i) connecting the current method with the backbone initialized with other remote sensing pre-training methodologies like [6, 7] rather than ImageNet-based one; (ii) extending this work to other remote sensing understanding tasks such as segmentation and change detection; (iii) designing novel freezing schedulers for further improvement. We believe these research directions are crucial for future work.

### v4

@@ 6 Conclusions and future work
In this paper, we have proposed a simple but efficient backbone fine-tuning strategy named DBF for the remote sensing object detection. This work demonstrated that alternating between freezing and unfreezing the backbone aids remote sensing object detection models in learning both general and domain-specific knowledge under long-term training. Particularly with the proposed method, a significant decrease in memory usage and training time can be achieved under very long training schedules. Our approach has shown powerful results on DOTA and DIOR-R datasets with various detection architectures. Additionally, the intuitive design of DBF guarantees high compatibility with current deep learning frameworks, facilitating easy comprehension and seamless integration.
@@
In the future, more sophisticated studies will be conducted as follows: (i) connecting the current method with the backbone initialized with other remote sensing pre-training methodologies like [19, 20] rather than ImageNet-based one; (ii) extending this work to other remote sensing understanding tasks such as segmentation and change detection; (iii) designing novel freezing schedulers to mitigate class imbalance issues [70] under long-term training. We believe these research directions are crucial for future work.
[…]
Author contributions JP and YK: Data curation, conceptualization, methodology and experimenting. MJ: Methodology and supervision.
[…]
Funding
This work was supported by the Agency For Defense Development Grant funded by the Korean Government (UI220066WD).
[…]
Availability of data and materials
Not applicable.
[…]
Code availability
Our code will be released on GitHub after our paper is accepted.
Declarations
[…]
Ethics approval and consent to participate Not applicable.
[…]
Consent for publication Not applicable.
[…]
Competing interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.